ROTATIONS OF CO_{2} AND CH_{4} ISOTOPOMERS IN HELIUM DROPLETS

\underline{DMITRY S. SKVORTSOV}, Chemistry Department, University of Southern California, Los Angeles, California 90089; HIROMICHI HOSHINA, Terahertz Sensing and Imaging Laboratory, RIKEN; 519-1399 Aramaki Aoba, Aoba-ku, Sendai, Miyagi, 980-0845, Japan; BORIS G. SARTAKOV, General Physics Institute, Vavilov str. 38, 117942 Moscow, Russia; and ANDREY F. VILESOV, Chemistry Department, University of Southern California, Los Angeles, California 90089.

The effect of isotopic substitution on molecular rotations of CO_{2} and CH_{4} have been studied via infrared spectroscopy in helium droplets. In the spectral region of ν_{1} +ν_{3} (3500 - 3700 cm^{−1}), rovibrational spectra of C^{18}O_{2}, C^{16}O_{2} and C^{16}O^{18}O have been obtained showing significant participation of surrounding helium in rotations. In contrast, minor changes of rotational constants of CH_{4} and CD_{4} have been derived from the rovibrational spectra in the ν_{3} region (3020 and 2260 cm^{−1} respectively). Rotational constants and band origins for all species have been determined. Coupling of molecular rotations with helium will be discussed.